OpenStudy (anonymous):

Question 3 (Multiple Choice Worth 1 points) Solve for x: 5x + 2 = 4x − 9 7 −7 11 −11 -------------------------------------------------------------------------------- Question 4 (Multiple Choice Worth 2 points) Solve for x: 3(x + 1) = −2(x − 1) + 6 1 4 5 25

OpenStudy (anonymous):

I guess Albert is solving the first one... so Here's the second: 3(x+1)=-2(x-1)+6 3x+3=-2x+2+6 3x+2x=2+6-3 ...

OpenStudy (albert0898):

[Question 3] 5x + 2 = 4x - 9 *Add nine to both sides 5x + 11 = 4x *Subtract 5x from both sides 11 = -1x *Divide both sides by -1 -11 = x